# DIPLODOCUS I: Framework for the evaluation of relativistic transport equations with continuous forcing and discrete particle interactions

### v2

@@ Abstract
DIPLODOCUS (Distribution-In-PLateaux methODOlogy for the CompUtation of transport equationS) is a novel framework being developed for the mesoscopic modelling of astrophysical systems via the transport of particle distribution functions through the seven dimensions of phase space, including continuous forces and discrete interactions between particles. This first paper in a series provides an overview of the analytical framework behind the model, consisting of an integral formulation of the relativistic transport equations (Boltzmann equations) and a discretisation procedure for the particle distribution function (Distribution-In-Plateaux). The latter allows for the evaluation of anisotropic interactions, and generates a conservative numerical scheme for a distribution function’s transport through phase space.
I Introduction
@@
The physical properties of astronomical phenomena must usually be inferred by distant observation alone. Blazars are a class of objects whose observed emissions are thought to originate from magnetised, relativistic jets of material formed close to the black holes at the centre of Active Galactic Nuclei (AGN) and directed almost exactly along our line of sight. The spectra of such objects is truly multi-wavelength, spanning from radio to ultra-high energy gamma rays, typically with a double peaked shape [1, 2, 3].
@@
The processes that dominate this emission are generally thought to be synchrotron radiation and inverse Compton scattering from either leptonic and/or hadronic populations within the jets. The observed brightness of such processes is heavily dependent on the properties of the in-situ particle populations, therefore it is of critical importance that such processes and populations are well modelled.
Modelling of AGN jets can broadly be classified as microscopic or macroscopic. These are not completely separate as the latter is derived from the former, however their execution is sufficiently distinguishable that they may be considered separate approaches.
@@
In deriving macroscopic (hydrodynamic) models such as (General Relativistic Magneto-)HydroDynamics (GRM)HD, particle populations are assumed to be follow some Maxwellian (thermal) distribution in momentum space, which depend on a small set of parameters, e.g. temperature, density and bulk flow velocity. This reduces the dimensionality, allowing a large spatial domains to be modelled, leading to high resolution simulations of the global jetted AGN structure [4, 5, 6, 7] at the cost of limited information on the underlying micro-scale interactions between particle populations, hence poor direct simulation of emissions.
@@
Microscopic (kinetic) models maintain information about the underlying distribution of particles in momentum space, at the cost of the increased dimensionality. To make this computationally tractable, other assumptions, e.g. isotropy in momentum, are often made to reduce the dimensionality as much as possible.
@@
One frequently used microscopic approach, in the field of plasma physics, is that of Particle-In-Cell (PIC). PIC replaces particle distributions with a finite number of discrete particles that attempt to simulate the entire distribution. This approach works well when distributions are near thermal or when small scale plasma instabilities and dynamics require resolution, such as for particle acceleration in magnetised shock fronts and reconnection regions [8, 9, 10, 11]. However, due computational limits, PIC can only simulate a finite number of particles. This limits sampling of both the distribution and interaction processes and constrains the spatial extent of any simulation. Therefore, PIC cannot typically be applied on a macro-scale.
@@
An alternate microscopic approach is to assume symmetries in the particle distribution in space and momentum space. For example, an emitting region of an astrophysical source can be simplified to a spherical or slab-like zone or zones with an internally isotropic distribution of particles. Early astrophysical use of these methods was in modelling cosmic ray transport [12, 13, 14]. Here, in addition to spherical symmetry in momentum space, specifically with reference to a co-moving frame, particle distributions were taken to be close to Maxwellian. The assumption of isotropy in particular led to extensive work [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] describing a wide range of interactions between particles and fields, with only small ventures into the world of anisotropy for specific [26, 27, 28, 29] and general [30] interactions. These works form the foundation of modern single/multi-zone models for jetted astrophysical sources [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], which maintain the tradition of momentum-space isotropy to limit computational requirements. These requirements are further reduced by considering additional symmetries in the spatial distribution of particles. In general, such models operate by considering independent, homogenous zones of isotropic plasma, normally spherical or slab-like in shape, with prescribed internal properties such as bulk Lorentz factor, magnetic field strength and initial particle distributions. These zones may then placed along the axis of the jet to generate is structure, though typically zones do not interact with one another. Within these zones, the particle distributions are evolved according to their interactions and the prescribed parameters of the zone. Emission spectra are then generated by combining and Doppler boosting photon populations from each zone into an observer frame. These spectra can then be fit to observed data of astrophysical sources. These zoned model fits have provided great insight into the environment and dynamics of sources like AGN jets, allowing the study of questions such as how are particles accelerated to the initial distributions, and whether the jet material dominated by leptonic or hadronic particles, which are still of some debate [46, 47, 48].
@@
The insight these zoned models have provided into blazars, and jetted astrophysical sources in general, such as X-Ray Binaries (XRB) and Gamma Ray Bursts (GRBs), should not be understated, see, for example, Refs. [49, 34, 50, 51, 38, 52, 53]. However, zoned models are intrinsically limited by their constraining symmetries and it is yet to be seen if their results are robust upon relaxation of these assumptions.
The work described in this series of papers was spawned in an attempt to assess that robustness. If momentum-space isotropy and spatial homogeneity were eliminated from zoned models, would the simulated emissions continue to fit the observed data? If so would their be a significant deviation in terms of parametrisation of the source? In either case, would this relaxation allow for deeper understanding?
Recent work has begun to question and probe these assumptions. Studies in the fields of pair-plasma winds [54, 55, 56] and cosmic-ray propagation [57, 58] have demonstrated self-consistent evolution of axisymmetric particle distributions using microscopic approaches. This shows that the additional dimensionality associated with relaxing the assumption of isotropy is within the capability of modern computing.
@@
Further, the axisymmetry studied by Ref. [58] provided a unique explanation for observations of a globular cluster, Terzan 5. A source of low-energy gamma rays is observed coincident with the location of the globular cluster, while a source of high-energy gamma rays is spatially displaced in the direction of Terzan 5’s magnetotail. Accelerated cosmic rays escape down this magnetotail in a narrow cone, with pitch angles not within our line of sight, so their high-energy emission is initially invisible to us. As they propagate down the magnetotail, their pitch angles begin to isotropise due to magnetic turbulence and cooling. Once sufficiently isotropic, their emissions become visible. This explanation of the observed spatial separation is only made possible by considering the evolution of both isotropic and anisotropic populations.
@@
Without the assumption of symmetries, kinetic models boil down to a question of particle transport: how does a particle go from point $A$ to point $B$? (With the slight complication that point $A$ and point $B$ exist not in three dimensions of space but the seven dimensions of phase space that include space, time and momentum.) This question expands beyond that of just modelling jetted astrophysical sources, and as such the range of fields this work may be applicable to is greatly expanded.
@@
This process of particle transport includes aspects of both microscopic and macroscopic models; small scale interactions and large scale advection and spatial structure. As such it lies between these two - in the mesoscopic regime.
This first paper (Paper I) in the series aims to give an overarching description of this mesoscopic particle transport in a very general sense, forming the backbone of this work. Section˜II describes, in a coordinate-free way, how particles get from $A$ to $B$ in phase space, and what exactly that means. Section˜III then adds in coordinates to the description. Finally Section˜IV describes the methodology this work is using to solve this problem of transport. Paper II, will then describe the numerical implementation of the transport defined in Paper I, with a demonstration of DIPLODOCUS’ microscopic capabilities, i.e. particle, interactions, emission and forcing. Paper III will then examine DIPLODOCUS’ macroscopic capabilities, with the inclusion of spatial structure and advection.
II Coordinate-Free Transport Equation
@@
This section defines (with some rigour) a coordinate-free, covariant form of the transport equations for particle distribution functions through phase space, also more generally known as Boltzmann equations [61]. The use of a coordinate-free approach is mathematically compact, and allows the transport equations to be adapted to any physical (or even non-physical) systems.
@@
There is a rich history of developing Boltzmann equations for the purpose of astrophysical modelling. However, most derivations focus solely on conservative cases, neglecting non-conservative elements. This work primarily builds on the approaches taken by Refs. [12, 62, 63]. Other works, some derived from the above sources and some independent, that have made impressions on the work presented here include Refs. [64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74].
II.1 A to B: Phase Flow on a Manifold
As a first step, consider a single particle located at some point $A$ at some instance and at $B$ at a later instance. These points $A$ and $B$ exists in an abstract phase space, a differentiable manifold of dimension $n$, denoted by $\mathcal{M}$. Both points exist along an unique curve $\gamma(\lambda)$, where $\lambda$ is an affine parameter along that curve such that $\gamma(\lambda_{A})=A$ and $\gamma(\lambda_{B})=B$. The coordinates of any point along this curve have the form $\bm{y}=\left(y^{1},...,y^{n}\right)$ and there exists a unique vector $\bm{v}$ tangent to the curve at each point. The curve $\gamma(\lambda)$ is therefore defined by the equation of motion
@@ II.1 A to B: Phase Flow on a Manifold
and is known as the particle’s worldline through phase space (see Fig.˜1).
@@
Now consider expanding this concept to every point $q\in\mathcal{M}$, i.e. at every point there exists a vector field $\bm{v}$ which defines a unique congruence of worldlines $\gamma_{q}(\lambda)$ that fill $\mathcal{M}$. These curves define how any particle will evolve if placed at any point $q\in\mathcal{M}$ (see Fig.˜2).
@@
II.2 The Distribution Function
For Hamiltonian (conservative) systems, $\bm{\mathrm{d}}\bm{\omega}=0$ and therefore $\mathcal{L}_{\bm{v}}\bm{\Omega}=0$, which is a statement of Liouville’s theorem [79] that the phase space volume is invariant with respect to the phase flow. Systems of importance to astrophysics are not in general Hamiltonian (see Section˜III.1) and as such the phase space volume is not, in general, invariant with respect to the phase flow i.e. $\mathcal{L}_{\bm{v}}\bm{\Omega}\neq 0$. Hence Liouville’s theorem does not apply and instead must be generalised.
@@ II.2 The Distribution Function
For non-trivial $\bm{\Omega}$, Eq.˜6 provides a differential transport equation for $f$ through phase space
@@
where $\partial Q$ is the bounding surface of $Q$. The measure $\int_{\partial Q}f\bm{\omega}$ can then freely be identified with the number of worldlines (particles) passing through the closed hypersurface $\partial Q$, and $f\bm{\omega}$ the flux of particles through $\partial Q$. Hence Eq.˜8 simply states that the divergence in phase space of the particle flux is equal to zero, i.e. the number of worldlines (particles) entering the region $Q$ is equal to the number leaving (see Fig.˜3). Therefore $f$ may be defined as the phase space density of worldlines (particles) on hypersurfaces in $\mathcal{M}$ or, more colloquially, the particle distribution function.
@@
II.3 Adding Collisions
The differential and integral transport equations for the distribution function, Eqs.˜7 and 8, assume an absence of discrete, discontinuous interactions (collisions) within $\mathcal{M}$ between particles. A collision at a point $q\in\mathcal{M}$ may be defined as a termination or starting of a worldline (more precisely the occupation of a worldline by a particle) due to a particle being destroyed/created or receiving a discrete transfer of momentum. For example, a particle $\mathfrak{a}$ [^3] Particle species are referred to by Gothic scripts in this paper. at a position $q_{\mathfrak{a}}$ in its phase space $\mathcal{M}_{\mathfrak{a}}$ may receive a discrete quantity of momentum due to scattering off a particle $\mathfrak{b}$ at position $q_{\mathfrak{b}}\in\mathcal{M}_{\mathfrak{b}}$. The scattering causes the termination of the worldlines of particles $\mathfrak{a}$ and $\mathfrak{b}$ at points $q_{\mathfrak{a}}$ and $q_{\mathfrak{b}}$ respectively as each particle’s momentum (a coordinate in phase space) has changed by a discrete amount. The worldlines for particles $\mathfrak{a}$ and $\mathfrak{b}$ are then re-started at some new positions $q^{\prime}_{\mathfrak{a}}$ and $q^{\prime}_{\mathfrak{b}}$ as shown in Fig.˜4. Alternatively, particle $\mathfrak{a}$ may undergo some emissive or decay process, where again it will undergo some discrete change in momentum and therefore terminate and re-start its worldline, while also starting new worldlines for any secondary particles emitted. Interactions, scattering, collisions, emissive and decay process all have the same effect in phase space, that being the termination and starting of worldlines, hence the term collision will used to refer collectively to these processes.
These collisions can then be represented by a new measure $\bm{C}(\bm{y})\propto\bm{\Omega}$ known as the collision integral, which “counts” the number of collisions within the volume $Q$ (see further in Section˜III.3). This measure can therefore be simply added to the right hand side of Eqs.˜7 and 8, to give the differential transport equations for $f$:
$\bm{v}(f)+f\text{div}_{\bm{\Omega}}\bm{v}=\bm{C}.$ (9)

### v4

# PaperIIPaper $\defcitealiasPaperIIIPaper II$I

@@ Abstract
DIPLODOCUS (Distribution-In-PLateaux methODOlogy for the CompUtation of transport equationS) is a framework being developed for the mesoscopic modelling of astrophysical systems via the transport of particle distribution functions through the seven dimensions of phase space, including continuous forces and discrete interactions between particles. This first paper in a series provides an overview of the analytical framework behind the model, consisting of an integral formulation of the relativistic transport equations (Boltzmann equations) and a discretisation procedure for the particle distribution function (Distribution-In-Plateaux). The latter allows for the evaluation of anisotropic interactions and generates a conservative numerical scheme for a distribution function’s transport through phase space.
1 Introduction
@@
Astronomical phenomena must usually be inferred by distant observation alone. Blazars are a class of objects whose observed emissions are thought to originate from magnetised, relativistic jets of material formed close to the black holes at the centre of Active Galactic Nuclei (AGN) and directed almost exactly along our line of sight. The spectra of such objects are truly multi-wavelength, spanning from radio to ultra-high-energy gamma rays, typically with a double-peaked shape (Urry and Padovani, 1995; Padovani et al., 2017; Blandford et al., 2019; Hovatta and Lindfors, 2019).
@@
The processes that dominate this emission are generally thought to be synchrotron radiation and inverse Compton scattering from either leptonic or hadronic populations, or a combination of both, within the jets. The observed brightness of such processes is heavily dependent on the properties of the in situ particle populations; it is therefore of critical importance that such processes and populations are well modelled.
@@
In deriving macroscopic (hydrodynamic) models such as (General Relativistic Magneto-)HydroDynamics (GRM)HD, particle populations are assumed to be highly collisional and in local thermal equilibrium. Therefore, knowledge of their distribution in momentum space may be neglected in favour of a small set of defining parameters, e.g. temperature, density and bulk flow velocity. This reduction of dimensionality permits large spatial domains to be modelled, leading to high resolution simulations of the global jetted AGN structure (McKinney and Gammie, 2004; McKinney, 2006; Chatterjee et al., 2019; Whitehead and Matthews, 2023) at the cost of limited information on the underlying micro-scale interactions between particle populations, hence poor direct simulation of emissions.
@@
By contrast, microscopic (kinetic) models maintain information about the underlying distribution of particles in momentum space, at the cost of the increased dimensionality. To make this computationally tractable, other assumptions, e.g. isotropy in momentum, are often made to reduce the dimensionality as much as possible.
@@
One frequently used microscopic approach, in the field of plasma physics, is that of Particle-In-Cell (PIC). PIC replaces particle distributions with a finite number of discrete particles that attempt to simulate the entire distribution. This approach works well when distributions are near thermal or when small scale plasma instabilities and dynamics require resolution, such as for particle acceleration in magnetised shock fronts and reconnection regions (Cerutti et al., 2013; Sironi et al., 2015; Mehlhaff et al., 2024; Comisso et al., 2024). However, due to computational limits, PIC can only simulate a finite number of particles. This limits sampling of both the distribution and interaction processes and constrains the spatial extent of any simulation. Therefore, PIC cannot typically be applied on a macroscopic scale.
@@
An alternative microscopic approach is to assume symmetries in the particle distribution in space and momentum space. For example, an emitting region of an astrophysical source can be simplified to a spherical or slab-like zone with an internally isotropic distribution of particles. Early astrophysical use of these methods was in modelling cosmic ray transport (Lindquist, 1966; Webb, 1985, 1989). Here, in addition to spherical symmetry in momentum space, specifically with reference to a co-moving frame, particle distributions were taken to be close to Maxwellian. The assumption of isotropy in particular led to extensive work (for example Blumenthal and Gould, 1970; Weaver, 1976; Aharonian and Atoyan, 1981; Svensson, 1982; Baring, 1987; Dermer, 1984; Brinkmann, 1984; Dermer, 1985, 1986; Coppi and Blandford, 1990; Sarkar et al., 2019) describing a wide range of interactions between particles and fields, with only small ventures into the world of anisotropy for specific (Stepney and Guilbert, 1983; Moskalenko and Strong, 2000; Kelner et al., 2014; Lai and Ng, 2023) and general (Everett and Cotter, 2024) interactions. These works form the foundation of modern single/multi-zone models for jetted astrophysical sources (Mastichiadis and Kirk, 1995; Mastichiadis et al., 2005; Dimitrakoudis et al., 2012; Potter and Cotter, 2012, 2013a, 2013b, 2013c; Böttcher et al., 2013; Cerruti et al., 2015; Gao et al., 2017; Potter, 2018; Fernández and van Eerten, 2021; Lucchini et al., 2022; Boula and Mastichiadis, 2022; Zacharias et al., 2022; Hahn et al., 2022; Klinger et al., 2024a; Stathopoulos et al., 2024; Cerruti et al., 2024), which maintain the tradition of momentum-space isotropy to limit computational requirements. These requirements are further reduced by considering additional symmetries in the spatial distribution of particles. In general, such models operate by considering independent, homogeneous zones of isotropic plasma, normally spherical or slab-like in shape, with prescribed internal properties such as bulk Lorentz factor, magnetic field strength and initial particle distributions. These zones may then be placed along the axis of the jet to generate its structure, though they typically do not interact with one another. The distributions of particles within each zone are then evolved according to their interactions and the parameters of the zone, with emission spectra generated by Doppler boosting escaping photons from each zone into an observer frame. These spectra can then be fit to observational data of astrophysical sources. These zoned model fits have provided great insight into the environment and dynamics of sources like AGN jets, allowing the study of questions such as how particles are accelerated to the initial distributions, and whether the jet material is dominated by leptonic or hadronic particles, questions remain a matter of debate (Weekes, 2003; Hoerbe et al., 2020; Kantzas et al., 2023).
@@
The insight these zoned models have provided into blazars, and jetted astrophysical sources in general, such as X-Ray Binaries (XRB) and Gamma-Ray Bursts (GRBs), should not be understated (see, for example Markoff et al., 2001; Böttcher et al., 2013; Marscher, 2014; Potter and Cotter, 2015; Potter, 2018; Kantzas et al., 2021; Klinger et al., 2024b). However, zoned models are intrinsically limited by their constraining symmetries and it is yet to be seen if their results are robust upon relaxation of these assumptions.
@@
Recent studies in the fields of pair-plasma winds (Aksenov et al., 2004, 2007, 2010) and cosmic-ray propagation (Krumholz et al., 2022, 2024) have demonstrated self-consistent evolution of axisymmetric particle distributions using microscopic approaches. Therefore, the additional dimensionality that comes with relaxing isotropy is within the capability of modern computing. Further, the study of the globular cluster Terzan 5 by Krumholz et al. (2024) provided a unique explanation for inconsistent gamma-ray emissions. Low-energy gamma rays are observed coincident with the location of the globular cluster, while a source of high-energy gamma rays is spatially displaced in the direction of Terzan 5’s magnetotail. This displacement was successfully modelled by the slow isotropisation of an accelerated population of cosmic rays propagating down the magnetotail. Initially, due to their narrow range of pitch angles, their emissions are beamed away from the line of sight and only once the distribution becomes sufficient isotropic do the emissions become visible. This explanation of the observed spatial separation was only made possible by considering the anisotropic evolution of the cosmic-ray population.
These findings raise many questions regarding the assumed momentum-space isotropy and spatial homogeneity of zoned models. If these assumptions were to be eliminated, would the simulated emissions continue to fit the observed data? Would this require a reinterpretation of the source, and could this provide deeper physical insights? These questions motivated the present work, which derives and details a framework designed to address them: DIPLODOCUS.
@@
With the elimination of assumed symmetries, kinetic models boil down to a question of particle transport: how does a particle go from point $A$ to point $B$? (With the slight complication that point $A$ and point $B$ exist not in three dimensions of space but the seven dimensions of phase space that include space, time and momentum.)
@@
This general approach to particle transport includes aspects of both microscopic and macroscopic models; small-scale interactions and large-scale structure. As such it lies between the two—in the mesoscopic regime. This regime, and the generality of the framework, permit its application to whole host of astrophysical phenomena, not just jetted sources.
[…]
2 Coordinate-Free Transport Equation
@@
This section defines (with some rigour) a coordinate-free, covariant form of the transport equations for particle distribution functions through phase space, also more generally known as Boltzmann equations (Boltzmann, 1872). The use of a coordinate-free approach is mathematically compact, and allows the transport equations to be adapted to any physical (or even non-physical) systems.
@@
There is a rich history of developing Boltzmann equations for the purpose of astrophysical modelling. However, most derivations focus solely on conservative cases, neglecting non-conservative elements. This work primarily builds on the approaches taken by Lindquist (1966); Ehlers (1971); Groot et al. (1980). Other works, some derived from the above sources and some independent, that have made impressions on the work presented here include Bichteler (1965); Hakim (1967a, b); Stewart (1971); Steeb (1979, 1980); Ellis et al. (1983a, b); Blandford and Eichler (1987); Cercignani (2002); Cardall and Mezzacappa (2003); Lifshits (2008); Shibata et al. (2014).
2.1 A to B: Phase Flow on a Manifold
As a first step, consider a single particle located at some point $A$ at some instance and at $B$ at a later instance. These points $A$ and $B$ exist in an abstract phase space, a differentiable manifold of dimension $n$, denoted by $\mathcal{M}$. Both points exist along a unique curve $\gamma(\lambda)$, where $\lambda$ is an affine parameter along that curve such that $\gamma(\lambda_{A})=A$ and $\gamma(\lambda_{B})=B$. The coordinates of any point along this curve have the form $\bm{y}=\left(y^{1},...,y^{n}\right)$ and there exists a unique vector $\bm{v}$ tangent to the curve at each point. The curve $\gamma(\lambda)$ is therefore defined by the equation of motion
@@ 2.1 A to B: Phase Flow on a Manifold
and is known as the particle’s worldline through phase space (see Fig. 1).
@@
Now consider expanding this concept to every point $q\in\mathcal{M}$, i.e. at every point there exists a vector field $\bm{v}$ which defines a unique congruence of worldlines $\gamma_{q}(\lambda)$ that fill $\mathcal{M}$. These curves define how any particle will evolve if placed at any point $q\in\mathcal{M}$ (see Fig. 2).
@@
2.2 The Distribution Function
For Hamiltonian (conservative) systems, $\bm{\mathrm{d}}\bm{\omega}=0$ and therefore $\mathcal{L}_{\bm{v}}\bm{\Omega}=0$, which is a statement of Liouville’s theorem (Liouville, 1838) that the phase space volume is invariant with respect to the phase flow. Systems of importance to astrophysics are not in general Hamiltonian (see Section 3.1) and as such the phase space volume is not, in general, invariant with respect to the phase flow i.e. $\mathcal{L}_{\bm{v}}\bm{\Omega}\neq 0$. Hence Liouville’s theorem does not apply and instead must be generalised.
@@ 2.2 The Distribution Function
For non-trivial $\bm{\Omega}$, Eq. (6) provides a differential transport equation for $f$ through phase space
@@
where $\partial Q$ is the bounding surface of $Q$. The measure $\int_{\partial Q}f\bm{\omega}$ can then freely be identified with the number of worldlines (particles) passing through the closed hypersurface $\partial Q$, and $f\bm{\omega}$ the flux of particles through $\partial Q$. Hence Eq. (8) simply states that the divergence in phase space of the particle flux is equal to zero, i.e. the number of worldlines (particles) entering the region $Q$ is equal to the number leaving (see Fig. 3). Therefore $f$ may be defined as the phase space density of worldlines (particles) on hypersurfaces in $\mathcal{M}$ or, more colloquially, the particle distribution function.
@@
2.3 Adding Collisions
The differential and integral transport equations for the distribution function, Eqs. (7) and (8), assume an absence of discrete, discontinuous interactions (collisions) within $\mathcal{M}$ between particles. A collision at a point $q\in\mathcal{M}$ may be defined as a termination or starting of a worldline (more precisely the occupation of a worldline by a particle) due to a particle being destroyed/created or receiving a discrete transfer of momentum. For example, a particle $\mathfrak{a}$ [^3] Particle species are referred to by Gothic scripts in this paper. at a position $q_{\mathfrak{a}}$ in its phase space $\mathcal{M}_{\mathfrak{a}}$ may receive a discrete quantity of momentum due to scattering off a particle $\mathfrak{b}$ at position $q_{\mathfrak{b}}\in\mathcal{M}_{\mathfrak{b}}$. The scattering causes the termination of the worldlines of particles $\mathfrak{a}$ and $\mathfrak{b}$ at points $q_{\mathfrak{a}}$ and $q_{\mathfrak{b}}$ respectively as each particle’s momentum (a coordinate in phase space) has changed by a discrete amount. The worldlines for particles $\mathfrak{a}$ and $\mathfrak{b}$ are then re-started at some new positions $q^{\prime}_{\mathfrak{a}}$ and $q^{\prime}_{\mathfrak{b}}$ as shown in Fig. 4. Alternatively, particle $\mathfrak{a}$ may undergo some emissive or decay process, where again it will undergo some discrete change in momentum and therefore terminate and re-start its worldline, while also starting new worldlines for any secondary particles emitted. Interactions, scattering, collisions, emissive and decay process all have the same effect in phase space, that being the termination and starting of worldlines, hence the term collision will be used to refer collectively to these processes.
These collisions can then be represented by a new measure $\bm{C}(\bm{y})\propto\bm{\Omega}$ known as the collision integral, which “counts” the number of collisions within the volume $Q$ (see further in Section 3.3). This measure can therefore be simply added to the right hand side of Eqs. (7) and (8), to give the differential transport equations for $f$:
$\bm{v}(f)+f\text{div}_{\bm{\Omega}}\bm{v}=\bm{C},$ (9)